OpenStudy (anonymous):

I am in Algebra 2 and need help with a question. does anybody think they could help??

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

6. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2. f(x) =6(3x+2) h(x)= 2x h(f(x))=

OpenStudy (anonymous):

if you cannot help that is totally okay :)

OpenStudy (anonymous):

Ohh... I see, you could simplify f(x) by distributing the 6, you get: f(x)=18x+12 to do the h(f(x)), you simply substitute f(x) to all x's in h(x) so, it'll be: h(f(x))=2(18x+12), then distribute the 2.

OpenStudy (anonymous):

oh thats easy. thank you so much. that helped alot. :)

OpenStudy (anonymous):

okay i need help with the second part now :/ how do i explain whether f(h(x)) and h(f(x)) will always result in the same number.????